OpenStudy (anonymous):

Solve: 3x − x + 2 = 12 −5 4 7

OpenStudy (anonymous):

3x − x + 2 = 12 3x - 1x = 2x So: 2x + 2 = 12 SUBTRACT THE 2 FROM BOTH SIDES 2x + 2 = 12 -2 -2 ---- ----- 2x = 10 DIVIDE BY 2 ON BOTH SIDES 2x = 10 --- --- 2 2 x = ____

OpenStudy (anonymous):

a

OpenStudy (anonymous):

x = 5

OpenStudy (anonymous):

dont know where you get the negative -5 must be your equation written wrong?

OpenStudy (anonymous):

ight i forgot to put d - 5 lol